OpenStudy (anonymous):

3(2x-4)=6(x-2) step by step plz

OpenStudy (anonymous):

Distribute 6x - 8 = 6x - 12 subtract 6x from both sides ... but when you do, the 6x's cancel each other out leaving -8 = -12 and since -8 does NOT = -12 the answer is NO SOLUTION

OpenStudy (agreene):

start by distributing. 3(2x-4)=6(x-2) 6x-12=6x-12 <===This is a true statement you can expand it to 0=0 but, either way there are infinite solutions @blexting cannot multiply correctly.

OpenStudy (anonymous):

6X-12=6X-12 BY DISTRIBUTIVE PROPERTY OF EQUALITY 0=0 ADDITION PROPERTY OF EQUALITY SO THEREFORE: NO SOLUTION

OpenStudy (agreene):

0=0 means there are infinite solutions \[x \exists \forall \mathbb{R}\]

OpenStudy (agreene):

For instance: x=9182374 3(2(9182374)-4)=6((9182374)-2) 55094232=55094232